OpenStudy (rahulmr):

\[3^{4\log_{3} (x-1)+2}\] can be simplified to (a) 9(x-1)^4 (b) 6(x-1)^4 (c) 2+(x-1)^4 (d) 9+(x-1)^4 (e) 4x-2 Help please. I don't know how should i simplify this. Thanks.

satellite73 (satellite73):

it can't be "simplified" to anything what does the actual question say?

OpenStudy (rahulmr):

That's all it says. I don't know what does it mean.

satellite73 (satellite73):

it doesn't mean anything there is no such mathematical operation as "simplify" for one thing for an other, all your possible answers seem to be exponential, not logs. perhaps it means "write in equivalent exponential form, but you would need an equation for that i

OpenStudy (rahulmr):

here, take a look at the image

satellite73 (satellite73):

oooh

satellite73 (satellite73):

you got a base of 3 there, didn't see it

OpenStudy (rahulmr):

yes

satellite73 (satellite73):

\[\huge 3^{4\log_3(x-1)+2}\]

satellite73 (satellite73):

now you can do something

satellite73 (satellite73):

first off, by one of the properties of the log, that 4 comes up as an exponent in \(\log_3(x-1)\) so as a first step turn \[\huge 3^{4\log_3(x-1)+2}\] in to \[\huge 3^{\log_3((x-1)^4)+2}\]

OpenStudy (rahulmr):

ok

satellite73 (satellite73):

then using another law of exponents, that \[b^xb^y=b^{x+y}\] you can make this \[\huge3^3\times 3^{\log_3((x-1)^4)}\]

satellite73 (satellite73):

oops should be \(3^2\) there, not \(3^3\)

OpenStudy (rahulmr):

i was going to say this

satellite73 (satellite73):

\[\huge9\times 3^{\log_3((x-1)^4)}\]

satellite73 (satellite73):

and since logs are the inverse of the exponential, you have \[3^{\log_3((x-1)^4)}=(x-1)^4\]

satellite73 (satellite73):

so final answer would be \[\huge 9(x-1)^4\]

OpenStudy (rahulmr):

ohh. now i get it. Thanks for the help. I was confused because of the 3.

satellite73 (satellite73):

yw